# Searching for Free-Floating Planets with TESS: I. Discovery of a First Terrestrial-Mass Candidate

### v1

@@ 4.4 Consistency with Expected Yield
Due to the expected increased abundance of FFPs with decreasing mass (Sumi et al., 2023b; Gould et al., 2022; Mroz & Poleski, 2023), events observed by TESS are expected to be largely due to FFPs with masses on the order of $10^{-2}-10^{-1}\,M_{\oplus}$, below which the typical FFP would not produce a detectable magnification given TESS’s sensitivity. As such, it is surprising to find an event consistent with a lens mass of $1-10\,M_{\oplus}$ in a search of a single sector. Rescaling Eq. (5) to $M=M_{\oplus}$, $N_{\star}=1.3\times 10^{6}$, $n=2.2~{}\text{pc}^{-3}$ (the best-fit value of Sumi et al. (2023b)), and $t_{\text{obs}}=23.6$ days (the average duration of Sector 61 observations after removal of poor quality data, and additional gap and edge cuts, §3.1), we find that the odds of detecting such an event are on the order of 1 in 100.
One possible explanation for our discovery, in light of low expected yields, is that we were simply lucky to have observed this candidate in our first TESS sector. Further searches of the TESS dataset may not uncover new candidates at the same rate of one event per sector. To test this assumption, we performed a preliminary search of an additional 6,213,765 QLP light curves across Sectors 62 – 65 following the same light curve processing, lensing search, and automated vetting procedure described here for Sector 61. While we have not yet manually reviewed or modeled all results to the same level of detail, we did not identify any new candidates similar to TIC-107150013. All similarly high-SNR candidates detected in these sectors were removed by the vetting procedure as asteroids or false positives with shapes inconsistent with a lens model. The only potential candidates remaining (to be explored in future works) possess significantly lower SNR than the event in Sector 61, and the signatures of a microlensing signal are not immediately apparent. If we revise our expected yield estimates from Eq. (5) with this larger dataset, we find that the probability of finding a terrestrial-mass lensing event within the data becomes approximately 1 in 20.
An alternative, more tantalizing explanation is that low-mass FFPs may be significantly more abundant than we expect. Given that TESS is uniquely suited to detecting a large number of terrestial-mass FFPs, we may only now be tapping into an enormous population of low-mass rogue worlds. Future searches of the TESS dataset will be crucial for better understanding the abundance and demographics of FFPs. Additionally, a more speculative explanation is that there exists a different population of dark lenses in this mass range, as has been suggested by previous studies of ground-based microlensing events (Mróz et al., 2017; Niikura et al., 2019). In particular, a macroscopic subcomponent of dark matter, such as primordial black holes (PBHs; see e.g. Carr & Kühnel (2022) for a review), fits our observation well. The high relative velocity we measure aligns with the expectation for a virialized halo of dark matter; furthermore, the detection of a single event in Sector 61 matches the abundance of the putative population of PBHs suggested by Niikura et al. (2019). [^6] For example, taking values within the parameter space suggested by Niikura et al. (2019) of $M=5\,M_{\oplus}$, $v_{\text{rel}}\approx 270$ km/s, and fraction of dark matter in PBHs $f_{\text{PBH}}=5\times 10^{-2}$, rescaling Eq. (5) yields an expected number of events in Sector 61 of $\approx 1$. If such a population of PBHs truly does exist, our ongoing search of other TESS sectors will provide key insight into its abundance and distribution.

### v2

@@ 4.2.3 Consistency with Expected Yield
Due to the expected increased abundance of FFPs with decreasing mass (Sumi et al., 2023; Gould et al., 2022; Mroz & Poleski, 2023), any events observed by TESS would be expected to be largely due to FFPs with masses on the order of $10^{-2}-10^{-1}\,M_{\oplus}$, below which the typical FFP would not produce a detectable magnification given TESS’s sensitivity. This is evident in Figure. 1. However, Figure. 1 also shows that in the entirety of existing observations, we expect only on the order of one (fiducial) to twenty ($1\sigma$) FFPs total.
Given this expected event rate, we can compute the odds of detecting at least one event in a search of five sectors. Summing over all bins in Fig. 1, we find that the central power-law predicts $\approx 1/70$ events per sector, averaged over sectors. Assuming detections follow a Poisson process, the probability of detecting at least one event in Sectors 61 – 65 is given by $1-P(k=0,\lambda\approx 1/14)\approx 7\%$, though this ranges from $\approx 0.7\%$ to $\approx 80\%$ for the $\pm 1\sigma$ slope of the mass function. These estimates indicate that the likelihood of detecting an FFP in the TESS data that we have searched is small. This motivates our investigation of alternatives to the lensing interpretation, as discussed in the following section.
4.3 False Positive Interpretation
Although TESS’s observing strategy allows for very high-cadence observation of target stars, its per-sector baseline is only 27.4 days. While many targets have been observed in multiple sectors, and hence have significantly longer baselines, these observations are sporadic and not conducive to searches for long-term variability. As such, comparison to external observations made by long-baseline stellar surveys is necessary for disentangling the nature of a signal. OGLE and ASAS-SN data (Shappee et al., 2014; Hart et al., 2023) described by Mróz (2023) revealed that the star associated with the Sector 61 signal exhibits a low-amplitude variability with a period of 424442-44 days, longer than the TESS baseline. Mróz (2023) suggested that this corresponds to rotational variability due to starspots, which in turn suggests the presence of magnetic activity in the target star. That work concluded that the observed event is likely a stellar flare.
In light of these observations, we analyze the characteristics of the primary event associated with TIC-107150013 as a stellar flare. As we show below, the apparent morphology of the event, its corresponding flare energy, and its relative isolation would make it a very unusual class of flare, and may point to a pernicious class of false positive that has not been well-characterized in the literature.
4.3.1 Consistency with Flare Morphology
Flares are expected to be largely asymmetric, with rapid flux increases followed by a slowly-falling tail, a morphology not exhibited in this particular event. We quantified this event’s consistency with flare morphology by fitting flare template models described by both Davenport et al. (2014) and Pitkin et al. (2014). The Davenport et al. (2014) model is commonly used to identify classical flares in dedicated studies of flares, and follows a fourth-order polynomial rise and double exponential decay characterized by the flare peak time ($t_{\mathrm{peak}}$), the peak amplitude ($A$), and the full time width at half the maximum flux ($t_{1/2}$). The Pitkin et al. (2014) model allows for a more flexible flare morphology including more symmetric curves, and involves a rapid rise with a half-Gaussian profile followed by an exponential decay, characterized by the flare peak time, the peak amplitude, the standard deviation of the Gaussian rise ($\tau_{g}$), and the exponential decay constant ($\tau_{e}$).
The fit results for all three models are provided in Table 3 and shown in Figure 12. We compare the BIC of these fits to the FSPL model and find $\mathrm{BIC}_{\mathrm{Pitkin}}-\mathrm{BIC}_{\mathrm{FSPL}}=12.8$ and $\mathrm{BIC}_{\mathrm{Davenport}}-\mathrm{BIC}_{\mathrm{FSPL}}=1262.9$. In other words, the Pitkin et al. (2014) model is the better flare model, but is still only $\exp{(-\Delta\mathrm{BIC}/2)}=0.0016$ times as probable as the FSPL model to minimize the information loss. Both models have essentially no support at the level of the data compared to the FSPL model. Furthermore, if we were to impose $\tau_{g}<\tau_{e}$ for the Pitkin et al. (2014) flare model, i.e., we require that the rise is shorter than the subsequent decay (following Pitkin et al. (2014) when using this model for flare detection), then model comparison gives $\mathrm{BIC}_{\mathrm{Pitkin}}-\mathrm{BIC}_{\mathrm{FSPL}}=98.9$ compared to the FSPL model. Therefore, the best-performing flare model becomes $3\times 10^{-22}$ times as probable as the FSPL model to minimize the information loss. These results indicate that existing models of flares struggle to account for events such as the one we have identified, and as such, this event provides a potential example of the new kinds of challenging backgrounds future space-based surveys will face
As mentioned in §4.2.1, there is a small secondary brightening spike during the rise of the event at $t=2986.85$ days, particularly visible in the residuals of the TGLC light curve; see left panel of Figure 12. We therefore also explore the false positive scenario that the primary event is not one flare, but the summation of two flares. We fit the sum of two Pitkin et al. (2014) flare models, and found $\mathrm{BIC}_{\mathrm{Double}}-\mathrm{BIC}_{\mathrm{FSPL}}=17.7$ ($64.8$ if we require $\tau_{g}<\tau_{e}$), which still indicates strong preference for the microlensing interpretation at the level of the data. We also fit the sum of an FSPL model and the Pitkin et al. (2014) flare model, and found $\mathrm{BIC}_{\mathrm{FSPL}}-\mathrm{BIC}_{\mathrm{FSPL+Flare}}=37.5$ ($35.3$ if we require $\tau_{g}<\tau_{e}$), which demonstrates that the combined FSPL and flare model gives the best fit overall, over FSPL alone, a single flare, and a double flare. However, because the FSPL fit parameters are within $1\sigma$ of the results assuming microlensing alone, we do not explore this scenario further in the paper.
While we prefer BIC over AIC because BIC better avoids overfitting for model selection (see §4.2.1), AIC is a common choice. When we used AIC for model comparisons we found that the FSPL model alone is still overwhelmingly preferred over both flare models ($\mathrm{AIC}_{\mathrm{Pitkin}}-\mathrm{AIC}_{\mathrm{FSPL}}=12.8$ and $\mathrm{AIC}_{\mathrm{Davenport}}-\mathrm{AIC}_{\mathrm{FSPL}}=1270.0$) as well as the double flare model when requiring $\tau_{g}<\tau_{e}$ ($\mathrm{AIC}_{\mathrm{Double}}-\mathrm{AIC}_{\mathrm{FSPL}}=36.4$). However, the double flare model is preferred over the FSPL model when relaxing this requirement ($\mathrm{AIC}_{\mathrm{FSPL}}-\mathrm{AIC}_{\mathrm{Double}}=10.6$). The discrepancy between BIC and AIC suggests that the double flare model is overfitting the data. We still find that the combined FSPL and flare model is the overall best performing model according to AIC ($\mathrm{AIC}_{\mathrm{FSPL}}-\mathrm{AIC}_{\mathrm{FSPL+Flare}}=65.9$), including over the double flare model. Our conclusion that the event is best explained by microlensing at the level of the data remains unchanged.
More generally, to our knowledge, no dedicated study of the rate of microlensing-like symmetric flares has been performed. We attempted to estimate the frequency of such events from the catalog of flares compiled by Günther et al. (2020a) from TESS Sectors 1 and 2 taken at 120-second cadence, as this provides the most direct comparison to our data. We visually inspected 291 flares with properties most similar to microlensing events, based on their timescales (full widths at half-maximum >0.03>0.03 days) and peak heights (amplitudes $0.5\%$), and flagged three events as symmetric corresponding to TIC-5640393, 32090583, and 271975726 (Figure 13). This could suggest a symmetric flare occurrence rate of $\sim 1\%$. However, our automated vetting pipeline failed all three events either due to high levels of red noise resulting in a low event SNR below our detection threshold, or other strong flares in the light curve resulting in the event being considered non-unique, and FSPL fits to these events failed to converge. Nevertheless, using the maximum likelihood results after 50,000 steps, we found that model comparison was unable to distinguish between interpretations to the same significance as TIC-107150013. TIC-5640393 corresponds to $\mathrm{BIC}_{\mathrm{Pitkin}}-\mathrm{BIC}_{\mathrm{FSPL}}=0.1$ while TIC-32090583 corresponds to $\mathrm{BIC}_{\mathrm{FSPL}}-\mathrm{BIC}_{\mathrm{Pitkin}}=0.1$, indicating no evidence for preferring either the flare or FSPL model ($\Delta\mathrm{BIC}<2$; Kass & Raftery, 1995). There is evidence for favoring the FSPL model over the best flare model for TIC-271975726 ($\mathrm{BIC}_{\mathrm{Pitkin}}-\mathrm{BIC}_{\mathrm{FSPL}}=2.6$), still significantly weaker evidence than what we found for TIC-107150013 ($\mathrm{BIC}_{\mathrm{Pitkin}}-\mathrm{BIC}_{\mathrm{FSPL}}=12.8$). Combined with the fact that these flares failed automated vetting and that their FSPL fits failed to converge, we conclude that the primary event associated with TIC-107150013 is unique compared to these other symmetric events based on morphology.
Table 3: Best-fit model results for the FSPL lensing model and Pitkin et al. (2014) and Davenport et al. (2014) flare models, for the primary event associated with TIC-107150013, a smaller bump earlier in the same light curve which could be a flare, and the three symmetric flares identified from the Günther et al. (2020a) flare catalog (§4.3.1). BIC values are included for model comparison. Because the FSPL fits only converged for the TIC-107150013 primary event and secondary bump. The FSPL properties for the symmetric flares were obtained after running emcee for 50,000 steps (2.5 million samples) which was not sufficient for convergence, and should therefore not be considered robust FSPL parameters.
Parameter TIC-107150013 (primary) TIC-107150013 (bump) TIC-5640393 TIC-32090583 TIC-271975726
FSPL lensing model
$t_{0}$ (BJD - 2457000) $2987.017_{-0.002}^{+0.002}$ $2972.224_{-0.013}^{+0.015}$ $1361.797_{-0.001}^{+0.001}$ $1366.014_{-0.001}^{+0.001}$ $1351.414_{-0.001}^{+0.001}$
$u_{0}$ $4.33_{-0.15}^{+0.15}$ $5.32_{-1.53}^{+2.78}$ $8.41_{-4.61}^{+0.70}$ $2.62_{-0.48}^{+1.23}$ $4.94_{-2.22}^{+2.01}$
$t_{E}$ (days) $074_{-0.002}^{+0.002} 0.0$$43_{-0.006}^{+0.006} 0.0$$05_{-0.001}^{+0.001} 0.0$$09_{-0.001}^{+0.001} 0.0$$007_{-0.002}^{+0.001}$
$\rho$ $4.57_{-0.15}^{+0.15}$ $4.78_{-3.25}^{+3.75}$ $8.73_{-6.02}^{+0.92}$ $1.9_{-1.3}^{+2.1}$ $5.14_{-3.54}^{+2.20}$
$f_{\mathrm{source}}~(e^{-}s^{-1})$ $875_{-52}^{+52}$ $865_{-53}^{+53}$ $150_{-7}^{+7}\times 10^{6}$ $701_{-59}^{+59}$ $942_{-56}^{+60}$
$f_{\mathrm{blend}}~(e^{-}s^{-1})$ $11_{-52}^{+52}$ $0_{-53}^{+53}$ $1_{-7}^{+7}\times 10^{6}$ $0_{-61}^{+57}$ $3_{-58}^{+58}$
BIC $56232.9$ $65048.4$ $758612.0$ $120883.6$ $166155.7$
Davenport et al. (2014) flare model
$t_{\mathrm{peak}}$ (BJD - 2457000) $2986.954_{-0.002}^{+0.002}$ $2972.164_{-0.006}^{0.005}$ $1361.790_{-0.001}^{+0.001}$ $1366.007_{-0.001}^{+0.001}$ $1351.407_{-0.001}^{+0.001}$
$t_{1/2}$ (days) $0.243_{-0.006}^{+0.006}$ $116_{-0.023}^{+0.017} 0.0$$13_{-0.002}^{+0.002} 0.0$$22_{-0.002}^{+0.001} 0.0$$019_{-0.002}^{+0.002}$
AA $069_{-0.004}^{+0.005} 0.0$$13_{-0.002}^{+0.002} 0.0$$24_{-0.003}^{+0.003} 0.0$$63_{-0.006}^{+0.007} 0.0$$035_{-0.003}^{+0.004}$
$f_{\mathrm{source}}~(e^{-}s^{-1})$ $860_{-55}^{+56}$ $860_{-54}^{+54}$ $150_{-7}^{+7}\times 10^{6}$ $694_{-61}^{+61}$ $937_{-61}^{+61}$
$f_{\mathrm{blend}}~(e^{-}s^{-1})$ $4_{-56}^{+55}$ $4_{-54}^{+54}$ $0_{-7}^{+7}\times 10^{6}$ $4_{-61}^{+61}$ $4_{-61}^{+61}$
BIC $57495.8$ $65035.5$ $758641.0$ $120943.4$ $166194.2$
Pitkin et al. (2014) flare model
$t_{\mathrm{peak}}$ (BJD - 2457000) $2987.081_{-0.005}^{+0.004}$ $2972.194_{-0.027}^{+0.042}$ $1361.805_{-0.002}^{+0.002}$ $1366.017_{-0.001}^{+0.001}$ $1351.418_{-0.002}^{+0.002}$
$\tau_{g}$ (days) $170_{-0.005}^{+0.005} 0.0$$67_{-0.03}^{+0.04} 0.0$$19_{-0.002}^{+0.002} 0.0$$15_{-0.002}^{+0.002} 0.0$$016_{-0.002}^{+0.002}$
$\tau_{e}$ (days) $0.091_{-0.004}^{+0.005}$ $165_{-0.031}^{+0.032} 0.0$$18_{-0.002}^{+0.003} 0.0$$16_{-0.002}^{+0.002} 0.0$$013_{-0.002}^{+0.003}$
AA $068_{-0.004}^{+0.005} 0.0$$7_{-0.05}^{+0.22} 0.0$$18_{-0.002}^{+0.002} 0.0$$58_{-0.006}^{+0.006} 0.0$$031_{-0.003}^{+0.003}$
$f_{\mathrm{source}}~(e^{-}s^{-1})$ $860_{-56}^{+55}$ $128_{-98}^{+413}$ $150_{-7}^{+7}\times 10^{6}$ $694_{-61}^{+62}$ $935_{-61}^{+60}$
$f_{\mathrm{blend}}~(e^{-}s^{-1})$ $4_{-55}^{+56}$ $736_{-413}^{+98}$ $0_{-7}^{+7}\times 10^{6}$ $4_{-61}^{+61}$ $6_{-60}^{+61}$
BIC $56245.8$ $65042.0$ $758612.1$ $120883.6$ $166158.4$
Figure 12: The primary brightening event associated with TIC-107150013 (left) and the secondary smaller bump earlier in the light curve (right) overlaid with colored lines showing the best-fit FSPL lensing model (black), Pitkin et al. (2014) flare model (orange), Davenport et al. (2014) flare model (green). We also show a Pitkin et al. (2014) flare + flare model (dotted orange) and FSPL + flare model (dotted black) for the primary event only (see text). The bottom panel shows the residual light curve and each model after removal of the baseline models chosen to interpret the events (FSPL model for the primary event; Davenport et al. (2014) model for the secondary event).
Figure 13: All flares manually flagged as symmetric and microlensing-like out of a subset of 291 flares compiled by Günther et al. (2020a). These flares would all have been rejected by our automated vetting tests due to large intrinsic red noise and the presence of other flares in the light curve, but suggest a possible symmetric flare occurrence rate up to $\sim 1\%$. See Sec. 4.3 for details.
Finally, there is a smaller, flare-like brightening feature in the light curve for TIC-107150013 near 2972 days (Figure 12), which could indicate potential activity on the star. Both flare models of this event are preferred over the FSPL model ($\mathrm{BIC}_{\mathrm{FSPL}}-\mathrm{BIC}_{\mathrm{Davenport}}=12.9$, $\mathrm{BIC}_{\mathrm{FSPL}}-\mathrm{BIC}_{\mathrm{Pitkin}}=6.4$; Table 2). However, this event has a low significance ($\mathrm{SNR}=6.5$), below our detection threshold, and fails our Uniqueness Test (§3.3.3) even after masking out the primary event at 2987 days. We also found that the light curve for TIC-1017150013 features significant amounts of red noise at the timescale of this event (0.32 days), where the amount of red noise ($\sigma_{r}=0.00096$) exceeds the amount of white noise ($\sigma_{w}/\sqrt{n}=0.00054$). A high amount of red noise is expected for high-precision light curves meant for transit searches (Pont et al., 2006). Combined with the low significance and lack of uniqueness, we cannot rule out this event as correlated noise.
4.3.2 Consistency with Flare Energies
We also computed the expected bolometric flare energy of the event, $E_{\mathrm{flare}}$, to assess its consistency with typical flares. Following the procedure described by Günther et al. (2020b), the luminosities of the star and flare in the TESS bandpass assuming black-body radiation are given by
$L_{\star}^{\prime}=\pi R_{\star}^{2}\int R_{\lambda}B_{\lambda}(T_{\mathrm{eff}})d\lambda,$ (8)
$L_{\mathrm{flare}}^{\prime}(t)=A_{\mathrm{flare}}(t)\int R_{\lambda}B_{\lambda}(T_{\mathrm{flare}})d\lambda,$ (9)
where $R_{\lambda}$ is the TESS response function, [^6] https://heasarc.gsfc.nasa.gov/docs/tess/data/tess-response-function-v2.0.csv $B_{\lambda}(T)$ is the Planck function evaluated at a given temperature $T$, and $A_{\mathrm{flare}}(t)$ is the time-dependent area of the flare. We solve for $A_{\mathrm{flare}}$ by recognizing that the normalized light curve gives the relative flare amplitude $(\Delta F/F)(t)=L_{\mathrm{flare}}^{\prime}(t)/L_{\star}$, giving
$A_{\mathrm{flare}}(t)=(\Delta F/F)(t)\pi R_{\star}^{2}\frac{\int R_{\lambda}B_{\lambda}(T_{\mathrm{eff}})d\lambda}{\int R_{\lambda}B_{\lambda}(T_{\mathrm{flare}})d\lambda}.$ (10)
We then find the bolometric flare luminosity $L_{\mathrm{flare}}(t)$ as
$L_{\mathrm{flare}}(t)=\sigma_{\mathrm{SB}}T_{\mathrm{flare}}^{4}A_{\mathrm{flare}}(t),$ (11)
and integrate over time to get the bolometric energy of the flare,
$E_{\mathrm{flare}}=\int L_{\mathrm{flare}}(t)dt.$ (12)
We take $T_{\mathrm{eff}}=4115$ K and $T_{\mathrm{flare}}=9000$ K for the star and flare, respectively, where 9000 K is a conservative lower limit consistent with previous studies (e.g. Shibayama et al., 2013; Günther et al., 2020a). We model $(\Delta F/F)(t)$ using the best-fit Pitkin et al. (2014) flare model. Integrating Eqn. 12 over the event, we find $E_{\mathrm{flare}}=1.3\times 10^{39}~\mathrm{erg}$, indicating that the event could constitute some form of unusually energetic superflare, which have typical energies of $10^{33}-10^{38}~\mathrm{erg}$ (e.g. Shibayama et al., 2013). The energy release returned by the above estimate is at least two orders of magnitude more energetic than all 8695 flares observed at 2 minute cadence in the first two sectors of TESS data (Günther et al., 2020a), the most energetic of which had $E_{\mathrm{flare}}=10^{36.9}~\mathrm{erg}$, which corresponded to the G-type giant star TIC-332487879. It would also be more energetic than all 57,961 flares with energies measured from TESS sectors 1 – 30 by Yang et al. (2023), of which only 10 (0.017%) had $E_{\mathrm{flare}}>10^{38}~\mathrm{erg}$ (Figure 14).
Figure 14: Histogram of flare energies for $57,961$ flares measured by Yang et al. (2023) from TESS sectors 1 – 30, compared to the measured flare energy of the primary brightening event associated with TIC-107150013.
The flare energies associated with the symmetric events around TIC-32090583 and 271975726 are also consistent with expectations for flares (Shibayama et al., 2013; Günther et al., 2020a; Yang et al., 2023), with $E_{\mathrm{flare}}=6.8\times 10^{34}~\mathrm{erg}$ and $4.7\times 10^{34}~\mathrm{erg}$, respectively. We cannot robustly estimate $E_{\mathrm{flare}}$ for TIC-5640393 because it lacks known stellar properties; however, assuming a Sun-like star we find $E_{\mathrm{flare}}=2.1\times 10^{35}~\mathrm{erg}$, and assuming a TIC-107150013-like star we find $E_{\mathrm{flare}}=3.6\times 10^{37}~\mathrm{erg}$, both of which are still in line with expectations for flare energies.
If the event identified on TIC-107150013 is interpreted as a flare, it is significantly more energetic than other observed flares and is poorly described by existing flare models. The event is unique compared to other symmetric events from Günther et al. (2020a) not only in terms of morphology, but also in terms of flare energy. While such a flare is possible, its lack of precedent in existing observations imply that it is very rare. We conclude that the flare false positive interpretation is unlikely, and, given the lack of an observational dataset of similar flares, we cannot quantitatively estimate how unlikely it is relative to the FFP interpretation.
4.3.3 Non-Flare False Positive Explanations
Other false positive scenarios are possible but are similarly challenged. For example, an event with a morphology similar to our signal has been observed on a young M-dwarf with an extremely strong (>1>1 KG) magnetosphere by (Palumbo et al., 2022). They interpret this event as the centrifugal breakout of a clump of stellar material that has been entrapped in a very strong rigidly rotating magnetosphere (RRM). Evidence for RRMs have been observed on hot A and B stars (Townsend et al., 2005; Oksala et al., 2012), and on very young M-dwarfs (Palumbo et al., 2022). However, RRMs require a strong (>1>1 KG) globally organized and stable magnetic field (Palumbo et al., 2022). The stellar properties of TIC-107150013 in Table 1 indicate that it is a cool, large, and therefore evolved, star which is not expected to have such a strong magnetosphere, hence this too remains an unlikely scenario.
TIC-107150013 appears to be a variable star based on OGLE data so, following common practice, we would not recommend TIC-107150013 to be considered in a general search for FFPs. Despite this, the signal we observe on TIC-107150013 provides an example of an FFP-like event arising from some rare astrophysical phenomenon that has not been thoroughly explored in the literature. Since TESS’s targets are nearby and their properties are well-characterized, our ongoing search through other TESS sectors will provide an opportunity to better understand such events and correlate them with the properties of the stars on which they appear, an important resource for future microlensing efforts.